y = 2sin¯ 1 3x

y = 2sin¯13x

Grade:Select Grade

1 Answers

Latika Leekha
9 years ago
Hello student,
As u haven’t mentioned what exactly is the question, I am assuming that you need the value of x.
We have y = 2sin-1 3x
So, y/2 = sin -1 3x
Hence, x = (1/3) sin (y/2).